What is the molar volume of an ideal gas if the pressure is 125 KPa at 300 K. If the experimental volume of the gas in a cube with 1.5 moles is 5 x 10-4 m3/mol, what is the compression factor?



1
Expert's answer
2021-01-22T06:23:14-0500

P= 125KPa , no. of moles(n) = 1.5 moles

T = 300K , V = ?

PV = n RT

125 × V = 1.5 × 0.082057 × 300

V = 0.2954L = 0.0002954m³/mol

compression factor = "\frac{5×10^-⁴}{0.0002954}"

= 1.69